# Experimentation on Endogenous Graphs

### v1

@@ 3 Models of Edge Formation
Just as estimating total treatment effects under exogenous interference requires imposing some additional structure on the interference mechanism (Basse and Airoldi, 2018), we will need to rule out arbitrary dependence of edges on treatment assignments. Let $S_{ar}$ be a subset of $\mathcal{R}$ satisfying $E_{ar}(\mathbf{T})=E_{ar}(\mathbf{T^{\prime}})$ for all $\mathbf{T},\mathbf{T}^{\prime}$ such that $T_{s}=T_{s}^{\prime}$ for all $s\in S_{ar}$. That is, whether the edge $(a,r)$ exists depends only on the treatment assignments to randomization units in $S_{ar}$. Non-exhaustive and non-mutually exclusive examples follow.
@@
  • r-driven edges. $S_{ar}=\{r\}$. Edges depend only on the randomization unit’s treatment status. This can be interpreted as a SUTVA assumption on edge formation: edge outcomes depend only on the treatment assignment to the randomization unit in question, not by treatment assignments to other units.
@@
For TTE estimation, we focus exclusively on the case of $r$-driven edges. This strikes a balance between meaningfully enriching the standard exogenous bipartite graph model, and still imposing enough structure on the problem to be able to derive useful statements on treatment effect magnitudes. This choice is also motivated by practicality—while setting $S_{ar}$ to be some strict superset of $\{r\}$ is more general, in applications there may be no clear principles available for constructing such sets.
@@ 4.1 Generalized Linear Exposure-Response Model
We specify the outcome model $y_{a}=\alpha_{a}+\beta_{a}x_{a}$, where $x_{a}=\sum_{r}T_{r}e_{ar}w_{ar}$ is the weighted number of treated edges $a$ is connected to in the realized, post-treatment graph, and $w_{ar}$’s are known weights. This outcome model satisfies the restrictions on potential outcomes stipulated in Definition 2, and is a generalization of the linear exposure-response model in Harshaw et al., (2021) to endogenous bipartite graphs. Outcomes are linear in the number of treated randomization units an analysis unit is exposed to, but the linear response is allowed to vary across analysis units. Treatments only affect outcomes through the observed edges, but the set of observed edges is potentially treatment dependent. By definition, the TTE is
@@ 4.2 Estimation Strategy
Considering $\beta_{a}$ first, Harshaw et al., (2021) propose $y_{a}(x_{a}-\mathbb{E}{x_{a}})/{\textrm{Var}(x_{a})}$ as an unbiased estimator in the exogenous bipartite graph case. This estimator is no longer feasible because the expectation and variance of $x_{a}$ are unknown, due to the endogenous nature of the graph edges. Moreover, as the examples in Section 2 show, naively applying this estimator assuming the realized graph is exogenous will lead to bias. To remedy this issue, we borrow from the literature on instrumental variables. Let $z_{a}^{u}=\sum_{r}T_{r}u_{ar}$ be a weighted exposure variable, where $u_{ar}$’s are fixed weights chosen by the researcher. If $\textrm{Cov}(x_{a},z^{u}_{a})$ is known and non-zero, we have a class of unbiased estimators for $\beta_{a}$,
@@ 4.3 Using Anchor Instruments
In general, even though the experimental design and hence the joint distribution of $\mathbf{T}$ is known, the covariance between the treatment intensity $x_{a}=\sum_{r}T_{r}e_{ar}w_{ar}$ and the instrument $z_{a}^{u}=\sum_{r}T_{r}u_{ar}$ is unknown. This is because it is a function of the unobserved values that the random variables $e_{ar}$ would take on under counterfactual treatment assignments. This suggests a path forward for a particular class of instruments for which $\textrm{Cov}(x_{a},z_{a}^{u})$ can be calculated. If there were a known set of pre-experiment edges which will continue to exist regardless of treatment assignments, this allows us to circumvent the difficulties posed by not observing $e_{ar}$ under counterfactual treatments. To start to formalize these observations, we present the definition of an anchor subgraph.
@@
An anchor subgraph is a set of edges that would exist under any treatment. [^4] For our Theorem 1 and Theorem 2, we can relax this assumption to requiring only $E_{ar}(\mathbf{1})=1$ under $r$-driven edges. Thus the existence of a known set of edges in the pre-experiment graph which satisfy the anchor subgraph property can be seen as a milder version of a “compliance”-type assumption familiar from the instrumental variables literature (Angrist et al., 1996), which in this setting would amount to requiring that all pre-treatment edges exist after the treatment has been applied. A subgraph in the pre-experiment data may plausibly satisfy the anchor subgraph property, in which case we can use it to construct an instrument for the endogenous $x_{a}$’s.
As a practical example, consider the case of Facebook Groups, where the bipartite graph consists of user-group memberships, with users being the analysis units and groups being the randomization units. Users outcomes are assumed to be a function of the treatments of groups to which they are connected, and those connections may be affected by the treatment in question. We may be able to identify a credible set of anchor-groups for a user, by selecting groups which they are highly unlikely to leave, based on a predictive model trained on pre-experimental data. Similarly, in the unipartite friendship graph, some long-lasting or active friendship connections may be highly likely to persist despite treatment assignment. These assumptions can be validated empirically—for any candidate anchor subgraph, we can estimate what fraction of edges disappear under treatment. To continue the Facebook groups example, less than $0.5\%$ of all group memberships disappear in a two-week window, and building models predictive of membership stability enables construction of a more stable anchor subgraph still. By contrast, experimental treatments may cause a larger number of new memberships to form over the same period, with a recent test generating a rate of edge formation of 0.75%.
@@
In the setting of Definition 2, suppose $G$ is a known anchor subgraph. If (a) edges are $r$-driven; (b) the treatment assignments are independent Bernoulli random variables with probability $p$; (c) $\{(a,r)\mid u_{ar}\neq 0\}\subset G$; (d) $c_{ar}=\mathbb{I}((a,r)\in G)$; (e) $w_{ar}\neq 0$ for all $(a,r)\in G$; and (f) |$|\{r\mid u_{ar}\neq 0\}|>0$ for all $a$, then
@@
To characterize the limiting behavior of the estimator proposed in Theorem 1, we consider an asymptotic regime in which the bipartite graph increases in size. If the graph is sufficiently sparse in this limit—i.e. the maximum degree of the analysis units and randomization units grows sufficiently slowly relative to the number of analysis units—then we can apply convergence results for dependent data, which allow us to derive consistency and asymptotic normality. Asymptotic normality relies on an application of Stein’s method (Stein, 1986; Ross, 2011), which has proven to be broadly applicable in demonstrating asymptotic normality in network settings (Chin, 2018; Harshaw et al., 2021). Theorem 2 below summarises these results. Assumptions (a), (b) and (c) ensure the instrument is positively correlated with the treatment exposure level $x_{a}$. Assumption (c) limits the total weight of randomization units connected to the same user and the discrepancy between those weights, and assumption (d) bounds outcomes.
Let $R_{a}=\{r\in\mathcal{R}\mid E_{ar}(\mathbf{1})=1\}$ be the set of randomization units connected to $a$ under treatment, and $A_{r}=\{a\in\mathcal{A}\mid E_{ar}(\mathbf{1})=1\}$ the set of analysis units connected to $r$ under treatment. Let $V_{a}=\{r\in\mathcal{R}\mid u_{ar}\neq 0\}$ be the set of randomization units connected to $a$ in the anchor subgraph.
@@
We maintain the assumptions in Theorem 1, and further assume (a) $|V_{a}|>0$; (b) $u_{ar}\geq 0$; (c) $W_{l}/|R_{a}|\leq|w_{ar}|\leq W_{h}/|R_{a}|$ for $r\in R_{a}$; and (d) $|y_{a}|\leq M$, where $p$, $1-p$, $W_{l}$, $W_{h}$, $M$ are constants bounded away from 00 and $\infty$. Then, letting $d_{\mathcal{A}}=\max_{a\in\mathcal{A}}|R_{a}|$ and $d_{\mathcal{R}}=\max_{r\in\mathcal{R}}|A_{r}|$,
  1. (i)
@@ 4.4 Statistical Testing under the Sharp Null
Based on estimators in (2), we can construct a statistical test for the sharp null $H_{0}^{\text{sharp}}$: $\beta_{a}=0$ for all $a$. Under $H_{0}^{\text{sharp}}$, $y_{a}=\alpha_{a}$ is independent of $\mathbf{T}$. We construct a modified version of (4)
@@
where we have made the dependence of $z_{a}^{u}$ on $\mathbf{T}$ clear. Under $H_{0}^{\text{sharp}}$ and the assumptions in Theorem 1, all randomness in $\tilde{\mu}^{u,c}$ comes from $z_{a}^{u}(\mathbf{T})$, the distribution of which is known.
@@
We maintain the assumptions in Theorem 2. Define the sharp null $H_{0}^{\textnormal{sharp}}$: $\beta_{a}=0$ for all $a$. The $H_{0}^{\textnormal{sharp}}$-null distribution of $|\tilde{\mu}^{u,c}|$ in (5) conditional on $\{y_{a}\mid a\in\mathcal{A}\}$ is known. Denote $c_{\alpha}$ as its $\alpha$-upper quantile, $\alpha\in(0,1)$. Rejecting if $|\tilde{\mu}^{u,c}|\geq c_{\alpha}$ is a valid size-$\alpha$ test for $H_{0}^{\textnormal{sharp}}$, and the power goes to $1$ if $d_{\mathcal{A}}^{3}d_{\mathcal{R}}/n_{a}\to 0$ and
lim $\liminf\left|\frac{1}{n_{a}}\sum_{a}\beta_{a}\sum_{r}(w_{ar}c_{ar})\right|>0.$

### v2

@@ 3 Models of Edge Formation
Just as estimating total treatment effects under exogenous interference requires imposing some additional structure on the interference mechanism (Basse and Airoldi 2018), we will need to rule out arbitrary dependence of edges on treatment assignments. Let $S_{ar}$ be a subset of $\mathcal{R}$ satisfying $E_{ar}(\mathbf{T})=E_{ar}(\mathbf{T^{\prime}})$ for all $\mathbf{T},\mathbf{T}^{\prime}$ such that $T_{s}=T_{s}^{\prime}$ for all $s\in S_{ar}$. That is, whether the edge $(a,r)$ exists depends only on the treatment assignments to the randomization units in $S_{ar}$. Examples of various edge formation assumptions follow.
@@
  • r-driven edges. $S_{ar}=\{r\}$. Edges depend only on the randomization unit’s treatment status. This can be interpreted as a SUTVA (Stable Unit Treatment Value Assumption) assumption on edge formation: edge outcomes depend only on the treatment assignment to the randomization unit in question, not by treatment assignments to other units.
@@
For TTE estimation, we focus exclusively on the case of $r$-driven edges. This strikes a balance between meaningfully enriching the standard exogenous bipartite graph model, and still imposing enough structure on the problem to be able to derive useful statements on treatment effect magnitudes. This choice is also motivated by practicality—while setting $S_{ar}$ to be some strict superset of $\{r\}$ is more general, in applications there may be no clear principles available for constructing such sets. Note that we are relaxing the common assumption of the interference graph being fixed, and in this respect generalize, rather than specialize, existing work—the $r$-driven edges assumption is trivially satisfied for a fixed graph. Some restrictions on edges and outcomes are indeed necessary to make meaningful claims about treatment effects in this more general setting.
We opt for the $r$-driven model of edge endogeneity as the simplest non-trivial generalization of exogenous edges. Robustness under the violation of this assumption and relaxing the assumption to $S_{r}$-driven edges, where edge formation can depend on peers’ treatments, are interesting future directions to explore. We wish to note that, while it is attractive to dispense with this assumption entirely, we see it as being analogous to the required restrictions on the interference structure in the exogenous graph case—without some structure, causal inference is impossible (see, e.g., Basse and Airoldi (2018)).
@@ 4.1 Generalized Linear Exposure-Response Model
We specify the outcome model $y_{a}=\alpha_{a}+\beta_{a}x_{a}$, where $x_{a}=\sum_{r}T_{r}e_{ar}w_{ar}$ is the weighted number of treated edges $a$ is connected to in the realized, post-treatment graph, and $w_{ar}$’s are known weights. The $w_{ar}$’s are a modeling choice, and we recommend the default choice of
$w_{ar}=1/(\text{\# pre-treatment edges connected to }a).$
The researcher might choose to incorporate domain knowledge into these weights based on how they expect a node to be affected by others (e.g., a user may be more affected by friends with whom they message frequently, or a buyer may be more affected by a seller they have frequently bought from in the past). These weights play the same role as the weights in the outcome model of Harshaw et al. (2021). This outcome model satisfies the restrictions on potential outcomes stipulated in Definition 2, and is a generalization of the linear exposure-response model in Harshaw et al. (2021) to endogenous bipartite graphs. Outcomes are linear in the number of treated randomization units the analysis unit is exposed to, but the linear response is allowed to vary across analysis units. Treatments only affect outcomes through the observed edges, but the set of observed edges is potentially treatment dependent. By definition, the TTE is
@@ 4.2 Estimation Strategy
Considering $\beta_{a}$ first, Harshaw et al. (2021) propose $y_{a}(x_{a}-\mathbb{E}{x_{a}})/{\textrm{Var}(x_{a})}$ as an unbiased estimator in the exogenous bipartite graph case. This estimator is no longer feasible because the expectation and variance of $x_{a}$ are unknown, due to the endogenous nature of the graph edges. Moreover, as the examples in Section 2 show, naively applying this estimator assuming the realized graph is exogenous will lead to bias. To remedy this issue, we borrow from the literature on instrumental variables. Let $z_{a}^{u}=\sum_{r}T_{r}u_{ar}$ be a weighted exposure variable, where $u_{ar}$’s are fixed weights chosen by the researcher. If $\textrm{Cov}(x_{a},z^{u}_{a})$ is known and non-zero, we have a class of unbiased estimators for $\beta_{a}$,
@@ 4.3 Using Anchor Instruments
In general, though the experimental design and hence the joint distribution of $\mathbf{T}$ is known, the covariance between the treatment intensity $x_{a}=\sum_{r}T_{r}e_{ar}w_{ar}$ and the instrument $z_{a}^{u}=\sum_{r}T_{r}u_{ar}$ is unknown. This is because it is a function of the unobserved values that the random variables $e_{ar}$ would take on under counterfactual treatment assignments. This suggests a path forward for a particular class of instruments for which $\textrm{Cov}(x_{a},z_{a}^{u})$ can be calculated. If there were a known set of pre-experiment edges that would continue to exist regardless of treatment assignments, this would allow us to circumvent the difficulties posed by not observing $e_{ar}$ under counterfactual treatments. To start formalizing these observations, we present the definition of an anchor subgraph.
@@
An anchor subgraph is a set of edges that would exist under any treatment. [^5] For our Theorem 1 and Theorem 2, we can relax this assumption to requiring only $E_{ar}(\mathbf{1})=1$ under $r$-driven edges. Thus the existence of a known set of edges in the pre-experiment graph which satisfy the anchor subgraph property can be seen as a milder version of a “compliance”-type assumption familiar from the instrumental variables literature (Angrist et al. 1996), which in this setting would amount to requiring that all pre-treatment edges exist after the treatment has been applied. A subgraph in the pre-experiment data may plausibly satisfy this property, in which case we can use it to construct an instrument for the endogenous $x_{a}$’s.
As a practical example, consider Facebook groups, where the bipartite graph links users (analysis units) to groups (randomization units). User outcomes depend on the treatment of connected groups, and those connections may themselves be affected by the treatment. To construct an anchor subgraph, we can identify groups a user is unlikely to leave—e.g., using a predictive model trained on pre-treatment data. Similarly, in unipartite settings, persistent friendships may serve as anchor edges. These assumptions can be validated empirically—for any candidate anchor subgraph, we can estimate what fraction of edges disappear under treatment. To continue the Facebook groups example, less than $0.5\%$ of all group memberships disappear over two weeks. By contrast, experimental treatments may cause a larger number of new memberships to form over the same period, with a recent test generating a rate of edge formation of 0.75%.
@@
In the setting of Definition 2, suppose $G$ is a known anchor subgraph. If (a) edges are $r$-driven; (b) the treatment assignments are independent Bernoulli random variables with probability $p$; (c) $\{(a,r)\mid u_{ar}\neq 0\}\subset G$; (d) $c_{ar}=\mathbb{I}((a,r)\in G)$; (e) $w_{ar}\neq 0$ for all $(a,r)\in G$; and (f) $|\{r\mid u_{ar}\neq 0\}|>0$ for all $a$, then
@@
Let $R_{a}=\{r\in\mathcal{R}\mid E_{ar}(\mathbf{1})=1\}$ be the set of randomization units connected to $a$ under treatment, and $A_{r}=\{a\in\mathcal{A}\mid E_{ar}(\mathbf{1})=1\}$ the set of analysis units connected to $r$ under treatment. Let $V_{a}=\{r\in\mathcal{R}\mid u_{ar}\neq 0\}$ be the set of randomization units connected to $a$ in the anchor subgraph. To see the time complexity of computing (4), we note that
$\hat{\beta}^{u}_{a}=\frac{y_{a}\sum_{r}u_{ar}(T_{r}-p)}{p(1-p)\sum_{r}w_{ar}u_{ar}}$
requires $\Theta(|V_{a}|)$ calculations, and $\widehat{\mathcal{W}_{a}(\mathbf{1})}^{c}$ requires $\Theta(|R_{a}|)$ calculations. Thus, the total time complexity is $\Theta\left(\sum_{a}|R_{a}|\right)$.
To characterize the limiting behavior of the estimator proposed in Theorem 1, we consider an asymptotic regime in which the bipartite graph increases in size. If the graph is sufficiently sparse in this limit—i.e. the maximum degree of the analysis units and randomization units grows sufficiently slowly relative to the number of analysis units—then we can apply convergence results for dependent data, which allow us to derive consistency and asymptotic normality. Asymptotic normality relies on an application of Stein’s method (Stein 1986, Ross 2011), which has proven to be broadly applicable in demonstrating asymptotic normality in network settings (Chin 2018, Harshaw et al. 2021). Theorem 2 below summarizes these results. Assumptions (a), (b) and (c) ensure the instrument is positively correlated with the treatment exposure level $x_{a}$. Assumption (c) limits the total weight of randomization units connected to the same user and the discrepancy between those weights, and assumption (d) bounds outcomes.
@@
We maintain the assumptions in Theorem 1, and further assume (a) $|V_{a}|>0$; (b) $u_{ar}\geq 0$; (c) $W_{l}/|R_{a}|\leq|w_{ar}|\leq W_{h}/|R_{a}|$ for $r\in R_{a}$; and (d) $|y_{a}|\leq M$, where $p$, 1p1-p, $W_{l}$, $W_{h}$, $M$ are constants bounded away from 0 and $\infty$. Then, letting $d_{\mathcal{A}}=\max_{a\in\mathcal{A}}|R_{a}|$ and $d_{\mathcal{R}}=\max_{r\in\mathcal{R}}|A_{r}|$,
  1. 1.
@@
With asymptotic normality and a variance estimator, we can construct confidence intervals. Section 4.4 provides a modified estimator with tractable variance that can be used as a proxy for uncertainty quantification for the original estimator. Section 5 shows this proxy variance yields confidence intervals with close-to-nominal coverage in simulations.
@@ 4.4 Statistical Testing under the Sharp Null
Based on the estimators of (2), we can construct a test for the sharp null $H_{0}^{\text{sharp}}$: $\beta_{a}=0$ for all $a$. Under $H_{0}^{\text{sharp}}$, $y_{a}=\alpha_{a}$ is independent of $\mathbf{T}$. Define a modified variant of (4),
@@
where the dependence of $z_{a}^{u}$ on $\mathbf{T}$ is explicit. Under $H_{0}^{\text{sharp}}$ and the assumptions of Theorem 1, all randomness in $\tilde{\mu}^{u,c}$ is from $z_{a}^{u}(\mathbf{T})$, the distribution of which is known.
@@
We maintain the assumptions in Theorem 2. Define the sharp null $H_{0}^{\textnormal{sharp}}$: $\beta_{a}=0$ for all $a$. The $H_{0}^{\textnormal{sharp}}$-null distribution of $|\tilde{\mu}^{u,c}|$ in (5) conditional on $\{y_{a}:a\in\mathcal{A}\}$ is known. Denote $c_{\alpha}$ as its $\alpha$-upper quantile, $\alpha\in(0,1)$. Rejecting if $|\tilde{\mu}^{u,c}|\geq c_{\alpha}$ is a valid size-$\alpha$ test for $H_{0}^{\textnormal{sharp}}$, and the power goes to $1$ if $d_{\mathcal{A}}^{3}d_{\mathcal{R}}/n_{a}\to 0$ and
$\liminf\left|\frac{1}{n_{a}}\sum_{a}\beta_{a}\sum_{r}(w_{ar}c_{ar})\right|>0.$
The time complexity to calculate (5) is $\Theta\left(\sum_{a}|V_{a}|\right)$. The exact $c_{\alpha}$ is expensive to compute, but noting that $\text{E}[\tilde{\mu}^{u,c}\mid\{y_{a}:a\in\mathcal{A}\},H_{0}^{\text{sharp}}]=0$, we can obtain a more conservative cutoff based on $\text{Var}(\tilde{\mu}^{u,c}\mid\{y_{a}:a\in\mathcal{A}\},H_{0}^{\text{sharp}})$, the time complexity to calculate which is also $\Theta\left(\sum_{a}|V_{a}|\right)$, because we have
$\text{Var}(\tilde{\mu}^{u,c}\mid\{y_{a}:a\in\mathcal{A}\},H_{0}^{\text{sharp}})=\frac{1}{p(1-p)}\sum_{r}\left(\sum_{a}\frac{y_{a}C_{a}u_{ar}}{n_{a}U_{a}}\right)^{2},$
where $C_{a}=\sum_{r}w_{ar}c_{ar}$ and $U_{a}=\sum_{r}w_{ar}u_{ar}$.
5 Simulations
5.1 Simulation Setting
We let $n_{a}=2000$ and $n_{r}=400$. We generate pre-treatment edges $e_{ar}^{\text{pre}}$’s as independent $\text{Bern}(0.1)$ random variables. Our algorithm takes the anchor edges to be the same as the pre-treatment edges. The edge outcome function is $r$-driven and defined as $E_{ar}(T_{r})=e_{ar}^{\text{pre}}+T_{r}(1-2e_{ar}^{\text{pre}})\text{Bern}(p_{\text{anchor}})$ if $(a,r)$ is anchor, and $e_{ar}^{\text{pre}}+T_{r}(1-2e_{ar}^{\text{pre}})\text{Bern}(p_{\text{non-anchor}})$ otherwise. That is, the edge outcome is the same as the pre-treatment edge if no treatment is applied to the randomization unit; otherwise, it will flip to the opposite of the pre-treatment edge with probability $p_{\text{anchor}}$ if the pre-treatment edge exists, and $p_{\text{non-anchor}}$ if it does not.
The $\alpha_{a}$’s are sampled from independent $\text{Unif}([0,1])$ variables, and the true treatment effect is set to be $\beta_{a}=\beta$ for all $a\in\mathcal{A}$. We run 1000 Monte Carlo simulations for various values of $\beta$, $p_{\text{anchor}}$ and $p_{\text{non-anchor}}$ and compare our anchor-based estimator (Anchor) with the ERL estimator (Harshaw et al. 2021) based on (a) pre-treatment graph ($\text{ERL}_{\text{pre}}$) and (b) post-treatment graph ($\text{ERL}_{\text{post}}$). The ERL estimators are
$\hat{\mu}^{\text{ERL}_{\text{pre}}}=\frac{1}{n_{a}}\sum_{a}\frac{y_{a}(x_{a}^{\text{pre}}(\mathbf{T})-p)}{p(1-p)/(\sum_{r\in\mathcal{R}}e_{ar}^{\text{pre}})}$
and
$\hat{\mu}^{\text{ERL}_{\text{post}}}=\frac{1}{n_{a}}\sum_{a}\frac{y_{a}(x_{a}^{\text{post}}(\mathbf{T})-p)}{p(1-p)/(\sum_{r\in\mathcal{R}}e_{ar}^{\text{post}})},$
where $x_{a}^{\text{pre}}(\mathbf{T})=\frac{\sum_{r\in\mathcal{R}}e_{ar}^{\text{pre}}T_{r}}{\sum_{r\in\mathcal{R}}e_{ar}^{\text{pre}}}$ and $x_{a}^{\text{post}}(\mathbf{T})=\frac{\sum_{r\in\mathcal{R}}e_{ar}^{\text{post}}T_{r}}{\sum_{r\in\mathcal{R}}e_{ar}^{\text{post}}}$. Here, $e_{ar}^{\text{pre}}$ and $e_{ar}^{\text{post}}$ are the pre-treatment and post-treatment edges between $a$ and $r$, respectively. We set $\beta=2$, $p_{\text{anchor}}=2\%$ and $p_{\text{non-anchor}}=0.2\%$ when they are not being varied.
5.2 Bias Comparison
Figure 2 shows the biases of the estimators under different parameter values. The anchor estimator results are consistent with unbiasedness with respect to different values of treatment effect ($\beta$) and edge formation intensity ($p_{\text{non-anchor}}$), as expected given our theoretical results. Using the post-treatment graph ($\text{ERL}_{\text{post}}$) naively can lead to severe overestimation due to the positive correlation between the marginal effect of an edge and the number of treated edges. Estimates based on the pre-treatment graph are still biased, but in the opposite direction, and the magnitude of the bias is much smaller as it stems solely from graph misspecification. Unbiasedness of the anchor estimator is achieved at the cost of higher variance, and the variance grows with $p_{\text{non-anchor}}$; with more edges being formed in response to treatment, we expect larger metric fluctuations between test and control contributed by $\widehat{\mathcal{W}_{ar}(\mathbf{1})}^{c}$, and hence, higher variance of the anchor estimator.
The anchor estimator is quite robust to violations of the anchor subgraph assumption, as we increase $p_{\text{anchor}}$. In Figure 2, we can see that its bias increases with $p_{\text{anchor}}$, but stays relatively small compared to the magnitude of the treatment effect, which is $2$. $\text{ERL}_{\text{pre}}$ is not directly affected by $p_{\text{anchor}}$ and is stable. On the other hand, the bias of $\text{ERL}_{\text{post}}$ changes its sign with the number of anchor edges disappearing. This is most likely due to the correlation between the marginal effect (which is positive) and the number of edges (that starts to diminish) becoming negative leading to underestimation of the true effect.
Figure 2: Bias with respect to different parameter values. Error bars are empirical standard deviations of the estimators over 1000 Monte Carlo runs.
Figure 3: 95%-CI coverage from 1000 independent Monte Carlo runs with respect to different parameter values. Error bars represent the $95\%$-CI for the true coverage probability.
5.3 Uncertainty Quantification
In order to provide a confidence interval (CI) for the anchor estimator, we appeal to asymptotic normality and use $\text{Var}(\tilde{\mu}^{u,c})$ under the sharp null as a proxy for the anchor estimator $\text{Var}(\hat{\mu}^{u,c})$, where $\tilde{\mu}^{u,c}$ is the estimator (5) defined in Section 4.4. We observe in Figure 3 that this approach delivers close to the nominal level coverage under different values of treatment effect ($\beta$) and anchor subgraph violation ($p_{\text{anchor}}$), with only modest under-coverage. By contrast, coverage is low when the edge formation density ($p_{\text{non-anchor}}$) is high, which is not surprising since the variance proxy is for an estimator that ignores the direct effect of newly formed edges.